7. Lori has the following utility function U = X0.5Y0.5 MUx = 0.5 X-0.5Y0.5 MUy = 0.5 X0.5Y-0.5

A.) Calculate Lori’s optimal consumption bundle when Px = Py = 10 given a budget of 200

B.)Calculate Lori’s optimal consumption bundle if Px = 5, other things equal

C.) Derive Lori’s demand for good X assuming it is linear.
Answer #1

Budget constraint be Pa + Peyzm 3 104+10y = 200 11x +y = 207-0 The Optimal Bundle Condition be | MRS = mue a P 1 muy y J ProGos xos yotas jo → 41 31275x) Put 24 = x Pn ego ③ we wel. > 5(24) + loy - 200 > y=10] *2.2 y 322110) >X220 The optimal Bund